OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −4

OpenStudy (anonymous):

Substitute the values in the expression: 4(-1) - 2(-4) = -4 + 8 = 4 is the answer...